We had been discussing if the extremely poor air quality due to the fire had effected our superoptics in the OFCs. To check this I performed a couple quick ringdown measurements on OFC1. For both wavelengths, the finesse of the cavity is close to measurements before the fires. It seems uneffected. For reference F = pi * c / L * (measured tau in screenshot).